Required information

Use the following information for the Exercises below.

[The following information applies to the questions displayed below.]

BMX Company has one employee. FICA Social Security taxes are 6.2% of the first $128,400 paid to its employee, and FICA Medicare taxes are 1.45% of gross pay. For BMX, its FUTA taxes are 0.6% and SUTA taxes are 5.4% of the first $7,000 paid to its employee.

Gross Pay through
August 31
Gross Pay for September
a. $ 6,100 $ 1,400
b. 2,800 2,900
c. 123,700 9,600

rev: 05_16_2019_QC_CS-168714

Exercise 9-5 Computing payroll taxes LO P2, P3

Compute BMX’s amounts for each of these four taxes as applied to the employee’s gross earnings for September under each of three separate situations (a), (b), and (c). (Round your answers to 2 decimal places.)

Required information Use the following information for the Exercises below. [The following information applies to the questioRequired information Use the following information for the Exercises below. [The following information applies to the questioExercise 9-6 Payroll-related journal entries LO P2 Assuming situation (a), prepare the employers September 30 journal entrieRequired information Use the following information for the Exercises below. [The following information applies to the questioUse the following information for the Exercises below. [The following information applies to the questions displayed below.)
Answer #1
1
a)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security 1400.00 6.20% 86.80
FICA-Medicare 1400.00 1.45% 20.30
FUTA 900.00 0.60% 5.40
SUTA 900.00 5.40% 48.60
b)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security 2900.00 6.20% 179.80
FICA-Medicare 2900.00 1.45% 42.05
FUTA 2900.00 0.60% 17.40
SUTA 2900.00 5.40% 156.60
c)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security 4700.00 6.20% 291.40
FICA-Medicare 9600.00 1.45% 139.20
FUTA 0 0.60% 0.00
SUTA 0 5.40% 0.00
Note: Earnings taxable for FUTA and SUTA in (a) = 7000-6100 = $900
Note: Earnings taxable for Social Security in (c) = 128400-123700 = $4700
2
Taxes to be Withheld From Gross Pay (Employee-Paid Taxes)
September Earnings Subject to Tax Tax Rate Tax Amount
Federal income tax 90.00
FICA—Social Security 1400.00 6.20% 86.80
FICA—Medicare 1400.00 1.45% 20.30
Total taxes withheld 197.10
Date General Journal Debit Credit
Sep 30 Salaries expense 1400.00
FICA—Social Security taxes payable 86.80
FICA—Medicare taxes payable 20.30
Employee federal income taxes payable 90.00
Salaries payable 1202.90
3
Employer Payroll taxes September earnings subject to tax Tax Rate Tax Amount
FICA-Social Security 1400.00 6.20% 86.80
FICA-Medicare 1400.00 1.45% 20.30
FUTA 900.00 0.60% 5.40
SUTA 900.00 5.40% 48.60
Total Payroll Taxes Expense 161.10
Date General Journal Debit Credit
Sep. 30 Payroll taxes expense 161.10
FICA—Social security taxes payable 86.80
FICA—Medicare taxes payable 20.30
Federal unemployment taxes payable 5.40
State unemployment taxes payable 48.60